OpenStudy (anonymous):

A ball is thrown across a playing field from a height of h = 6 ft above the ground at an angle of 45° to the horizontal at the speed of 20 ft/s. It can be deduced from physical principles that the path of the ball is modeled by the functiony = −32/((20)^2)x^2 + x + 6 where x is the distance in feet that the ball has traveled horizontally Find the maximum height attained by the ball. (Round your answer to three decimal places.) (b) Find the horizontal distance the ball has traveled when it hits the ground. (Round your answer to one decimal place.

OpenStudy (anonymous):

for max height differentiate...and put dy/dx = 0 find values of x..put em into the original eqn

OpenStudy (anonymous):

its not working for me can u plz try the solution for me

OpenStudy (anonymous):

20^2 in the denominator?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

dy/dx = -64x/400 + 1 = 0 x = 400/64 = 25/4

OpenStudy (anonymous):

now put in x=25/4 into ur original eqn

OpenStudy (anonymous):

yeah i kept the 6.25 in the solution but i am not getting the right answer

OpenStudy (anonymous):

9.125

OpenStudy (anonymous):

is it?

OpenStudy (anonymous):

IS IT?

OpenStudy (anonymous):

it works now

OpenStudy (anonymous):

I think I kept the wrong answer I got 15.375

OpenStudy (anonymous):

do u know the second part?

OpenStudy (anonymous):

well do calculations correctly

OpenStudy (anonymous):

what class r u in?

OpenStudy (anonymous):

college grade

OpenStudy (anonymous):

ohk now the second part

OpenStudy (anonymous):

when the ball falls down what is its height?

OpenStudy (anonymous):

o

OpenStudy (anonymous):

so set y = 0

OpenStudy (anonymous):

and find x

OpenStudy (anonymous):

wow u made it very easy for me

OpenStudy (anonymous):

dont sweat it

OpenStudy (anonymous):

wht were u thinking of?

OpenStudy (anonymous):

finding x value now

OpenStudy (anonymous):

found?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

iam unable to find

OpenStudy (anonymous):

use quadratic formula

OpenStudy (anonymous):

am using the same

OpenStudy (anonymous):

not getting

OpenStudy (anonymous):

am getting it as complex solution

OpenStudy (anonymous):

33.875

OpenStudy (anonymous):

how did u get it can u plz sho the steps when am using quadratic functions its giving me complex value

OpenStudy (anonymous):

33.875 is wrong answer@ him

OpenStudy (anonymous):

16.9375

OpenStudy (anonymous):

now right?

OpenStudy (anonymous):

how did u get that value can u plz show me the steps?

OpenStudy (anonymous):

\[x = (-b \pm \sqrt{b^2 - 4ac})/2a\]

OpenStudy (anonymous):

its exact value u got it perfect? I used the same formula but I got in complex

OpenStudy (anonymous):

a = -32/400 b= 1 c =6

OpenStudy (anonymous):

do it properly

OpenStudy (anonymous):

ull get it

OpenStudy (anonymous):

made a small mistake

OpenStudy (anonymous):

got it now

OpenStudy (anonymous):

made wrong coz of my time is running off thanx a lot for your help

OpenStudy (anonymous):

no problem click good answer

OpenStudy (anonymous):

If a ball is thrown directly upward with a velocity of 40 ft/s, its height (in feet) after t seconds is given by y = 40t − 16t2. What is the maximum height attained by the ball? (Round your answer to the nearest whole number.)

OpenStudy (anonymous):

16t^2 in the above one

OpenStudy (anonymous):

25 ft

OpenStudy (anonymous):

click on good answer on the right of my name

OpenStudy (anonymous):

how did u get it directly?

OpenStudy (anonymous):

i calculated in my head

OpenStudy (anonymous):

nothing direct about it

OpenStudy (anonymous):

did u click?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can u tell me the steps?

OpenStudy (anonymous):

h(t) = 40t - 16t^2 dh/dt = 40 -32t = 0 t = 1.25 put it in the eqn

OpenStudy (anonymous):

gimme a medal

OpenStudy (anonymous):

i already gave u

OpenStudy (anonymous):

doesnt show any..lol

OpenStudy (anonymous):

u became my fan

OpenStudy (anonymous):

i clicked on the medal is it not?

OpenStudy (anonymous):

click on "good answer" beside the medal

OpenStudy (anonymous):

njoi the day dude

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

u 2